# Smoothing Variances Across Time: Adaptive Stochastic Volatility

### v4

@@ 4.1 Set-up
DGP1: SV with 1 Regime DGP2: SV with 2 Regimes
$y_{t}\sim N(0,\exp(h_{t}))$ $h_{t}=3+0.8(h_{t-1}-3)+0.2u_{t}$ $u_{t}\stackrel{{\scriptstyle iid}}{{\sim}}N(0,1)$ $y_{t}\sim N(0,\exp(h_{t}))$ $h_{t}=m_{s_{t}}+0.8(h_{t-1}-m_{s_{t-1}})+0.2u_{t}$ $u_{t}\stackrel{{\scriptstyle iid}}{{\sim}}N(0,1)$ $m_{s_{t}}=\begin{cases}\,-10\,,&\text{if $s_{t}=0$.}\\ \,6\,,&\text{if $s_{t}=1$.}\end{cases}$
DGP3: SV with 3 Regimes DGP4: GARCH with 1 Regime
$y_{t}\sim N(0,\exp(h_{t}))$ $h_{t}=m_{s_{t}}+0.8(h_{t-1}-m_{s_{t-1}})+0.2u_{t}$ $u_{t}\stackrel{{\scriptstyle iid}}{{\sim}}N(0,1)$ $m_{s_{t}}=\begin{cases}\,-10\,,&\text{if $s_{t}=0$.}\\ \,-3\,,&\text{if $s_{t}=1$.}\\ \,3\,,&\text{if $s_{t}=2$.}\end{cases}$ $y_{t}\sim N(0,\sigma^{2}_{y,t})$ $\sigma_{y,t}^{2}=1+0.1y_{t-1}^{2}+0.5\sigma_{y,t-1}^{2}$
DGP5: GARCH with 2 Regimes DGP6: GARCH with 3 Regimes
$y_{t}\sim N(0,\sigma^{2}_{y,t})$ $\sigma_{y,t}^{2}=m_{s_{t}}+0.15y_{t-1}^{2}+\beta_{s_{t}}\sigma_{y,t-1}^{2}$ $m_{s_{t}}=\begin{cases}\,8\,,&\text{if $s_{t}=0$.}\\ \,0.1\,,&\text{if $s_{t}=1$.}\end{cases}$ $\beta_{s_{t}}=\begin{cases}\,0.8\,,&\text{if $s_{t}=0$.}\\ \,0.3\,,&\text{if $s_{t}=1$.}\end{cases}$ $y_{t}\sim N(0,\sigma^{2}_{y,t})$ $\sigma_{y,t}^{2}=m_{s_{t}}+0.15y_{t-1}^{2}+\beta_{s_{t}}\sigma_{y,t-1}^{2}$ $m_{s_{t}}=\begin{cases}\,12\,,&\text{if $s_{t}=0$.}\\ \,8\,,&\text{if $s_{t}=1$.}\\ \,0.1\,,&\text{if $s_{t}=2$.}\end{cases}$ $\beta_{s_{t}}=\begin{cases}\,0.8\,,&\text{if $s_{t}=0$.}\\ \,0.5\,,&\text{if $s_{t}=1$.}\\ \,0.2\,,&\text{if $s_{t}=2$.}\end{cases}$
DGP7: Sinusoidal DGP8: Piecewise Constant
$y_{t}\sim\mathcal{N}(0,\exp(h_{t}))$ $h=A\sin(10(2\pi t)/T)+B\cos(10(2\pi t)/T)$ $\quad+C\sin(3(2\pi t)/T)+D\cos(3(2\pi t)/T)$ $A,B,C,D\stackrel{{\scriptstyle ind}}{{\sim}}U(0,5)$ $y_{t}\sim\mathcal{N}(0,\exp(h_{t}))$ $j:=\left\lfloor\frac{t}{25}\right\rfloor+1$ $h_{t}=(-1)^{j}\cdot|z_{j}|$ $z_{j}\sim\begin{cases}N(5,0.5^{2})&\text{if $j$ is even}\\ N(0,0.5^{2})&\text{if $j$ is odd}\end{cases}$
Table 1: Summary of the eight data‐generating processes (DGP). DGP 1 - 3 simulate a stochastic‐volatility model with 1, 2, or 3 regimes; DGP 4–6 simulate a GARCH(1,1) model with 1, 2, or 3 regimes. In DGP 7, $h_{t}$ is a random sinusoid; in DGP 8, $h_{t}$ is piecewise constant with random intercepts. For DGPs 2 and 4, the transition probability of remaining in the same state is 0.98, and transitioning to a different state is 0.02. For DGPs 3 and 6, the probability of staying in the same state is 0.98, while switching to any other state occurs with probability 0.01. Each DGP comprises 1,000 sample paths, each path being 1,000 data points in length.
The simulation study evaluates the performance of the proposed ASV model against several existing methods, including the SV, Markov-Switching SV with two regimes (MSSV2), RWSV, GARCH, and two-regime GARCH (MSGARCH2). A local-only shrinkage model RWSV-BL is also included for comparison. Four variants of ASV: ASV with the horseshoe prior (ASV-HS), ASV with the dynamic horseshoe prior (ASV-DHS), and their respective versions with a nugget effect (ASV-HS-N and ASV-DHS-N) are considered. ASV-HS corresponds to the special case of Equation 4 with $a=b=1/2$ and $\phi=0$, while ASV-DHS generalizes this by allowing $\phi$ to be estimated from the data. The distinction between ASV with and without nugget is shown in Equation 5. All ASV models use $k=1$, applying shrinkage to the first-order differences of the latent log-variance process. This choice reflects the empirical observation that volatility in the SV or GARCH-type models typically exhibits clustering rather than gradual changes in slope.
Eight DGPs are considered: DGP 1–3 use SV models with one to three regimes; DGP 4–6 use GARCH models with one to three regimes; DGP 7 features sinusoidal log-variance with random coefficients; and DGP 8 involves piecewise constant log-variance with 3 to 5 random breakpoints. For each DGP, 1,000 sample paths are simulated. Detailed parameterization of the DGPs may be found in Table 1.
@@
$MAE$ $=\frac{1}{T}\sum_{t=1}^{T}|\sigma_{t}-\hat{\sigma_{t}}|,$ $EC$ $=\frac{1}{T}\sum_{t=1}^{T}\mathds{1}_{(\hat{\sigma}_{t,0.05},\hat{\sigma}_{t,0.95})}(\sigma_{t}),$ $MCIW$ $=\frac{1}{T}\sum_{t=1}^{T}(\hat{\sigma}_{t,0.95}-\hat{\sigma}_{t,0.05}).$
The MAE measures the accuracy of the volatility estimate by calculating the average absolute difference between the true volatility $\sigma_{t}$ and the model’s estimate $\hat{\sigma_{t}}$. For Frequentist methods, the point estimate is based on its maximum likelihood estimate, while it is the posterior mean for the Bayesian methods. The EC assesses the reliability of each model’s uncertainty quantification by calculating the proportion of the true volatility falling within the model’s 90% credible interval. Finally, the MCIW evaluates the precision of the uncertainty quantification, where a narrower credible interval represents a higher level of precision.
All models are implemented in R language (R Core Team, 2013). Specifically, the SV model is implemented via stochvol package (Kastner, 2016), MSSV2 as specified in Hamilton (1989) and RWSV are implemented by the authors as no readily available packages in R exist. GARCH(1,1) is fitted via fGarch package (Wuertz et al., 2023), and MSGARCH2 is fitted via MSGARCH package (Ardia et al., 2019). For fitting RWSV-BL, dsp package by Kowal et al. (2019) and genlasso package by Arnold and Tibshirani (2022) are used. dsp package by Kowal et al. (2019) is expanded by the authors to implement ASV. For Bayesian models, 5,000 posterior samples are generated after 20,000 burn-ins for each sample path.
@@ 4.2 Results
MAE SV 0.4183 (0.0195) 4.3355 (0.9853) 0.4948 (0.1645) 0.1641 (0.0212) 1.2965 (0.2978) 1.2804 (0.3159) 1.1831 (0.9398) 2.0796 (0.1713)
MSSV2 0.6288 (1.6236) 1.0265 (0.282) 0.4904 (0.1676) 0.3034 (1.1001) 2.6922 (2.9965) 3.1646 (3.4242) 2.6815 (2.5832) 2.019 (0.1632)
GARCH 0.4297 (0.0227) 5.8768 (0.817) 0.9623 (0.2776) 0.0421 (0.0181) 2.0226 (0.456) 1.8237 (0.437) 2.7572 (2.2338) 4.1509 (0.3525)
MSGARCH2 0.4334 (0.0318) 5.9983 (1.6226) 1.1078 (1.2224) 0.0577 (0.0206) 0.9559 (0.1991) 1.247 (0.26) 2.1427 (1.6771) 2.7224 (0.7748)
RWSV 0.4745 (0.0309) 4.2692 (0.9614) 0.4875 (0.161) 0.1704 (0.0199) 1.2744 (0.2905) 1.2576 (0.309) 1.1668 (0.925) 2.0301 (0.1692)
RWSV-BL 1.0103 (0.0597) 3.4705 (0.7682) 0.4641 (0.152) 0.2811 (0.0311) 1.4004 (0.3371) 1.4543 (0.3761) 1.42 (1.0283) 1.9933 (0.1658)
[…]
ASV-DHS 0.4517 (0.0313) 1.4892 (0.3192) 0.2369 (0.0725) 0.1777 (0.0212) 0.9462 (0.2124) 1.0336 (0.2535) 0.9426 (0.7442) 1.2209 (0.1385)
ASV-HS-N 0.4437 (0.0292) 1.4623 (0.3275) 0.236 (0.0746) 0.1776 (0.0209) 0.9754 (0.2228) 1.0617 (0.2588) 0.9358 (0.7373) 1.2449 (0.1386)
ASV-DHS-N 0.4477 (0.0304) 1.4892 (0.3205) 0.237 (0.0725) 0.1783 (0.0212) 0.9537 (0.2146) 1.0414 (0.2558) 0.9427 (0.7423) 1.2242 (0.1382)
MCIW SV 1.7151 (0.4996) 23.9914 (5.5341) 2.6681 (0.943) 0.8261 (0.0985) 6.9493 (1.6761) 6.9645 (1.7686) 6.9224 (5.4048) 10.5937 (0.8248)
MSSV2 1.6856 (0.8774) 6.1165 (2.2849) 2.9168 (0.9828) 1.2632 (1.4566) 8.1412 (1.9785) 8.4577 (2.134) 7.9837 (4.4592) 11.0375 (0.8436)
@@
RWSV 1.4682 (0.0678) 23.6249 (5.4109) 2.6127 (0.9215) 0.707 (0.1036) 6.7469 (1.6199) 6.7313 (1.7063) 6.801 (5.3228) 10.0901 (0.7914)
RWSV-BL 5.8851 (0.178) 19.7094 (4.3992) 2.5829 (0.8688) 1.6936 (0.1421) 7.9766 (1.9155) 8.479 (2.1574) 8.4838 (6.0041) 10.6267 (0.7472)
[…]
ASV-DHS 1.0711 (0.1338) 6.5254 (1.3109) 1.026 (0.3029) 0.6442 (0.1047) 4.0214 (0.8943) 3.9701 (0.9812) 4.3191 (3.3025) 6.3782 (0.4163)
ASV-HS-N 1.7522 (0.3342) 7.0255 (1.4975) 1.1211 (0.354) 0.7136 (0.1034) 4.4482 (1.0088) 4.4818 (1.0818) 4.6034 (3.4542) 6.6306 (0.4301)
ASV-DHS-N 1.7898 (0.3295) 7.1332 (1.431) 1.1215 (0.3332) 0.7127 (0.1017) 4.2736 (0.9445) 4.3395 (1.0468) 4.6343 (3.4943) 6.5572 (0.4303)
EC SV 0.8655 (0.0968) 0.9159 (0.0093) 0.9177 (0.0104) 0.9621 (0.0131) 0.9296 (0.0116) 0.9327 (0.0121) 0.9764 (0.0114) 0.9091 (0.0107)
MSSV2 0.8348 (0.1264) 0.9589 (0.0368) 0.939 (0.0113) 0.9839 (0.0176) 0.9451 (0.0105) 0.9514 (0.0103) 0.9757 (0.0085) 0.9173 (0.0101)
@@
RWSV 0.7779 (0.0276) 0.9144 (0.0094) 0.9148 (0.0107) 0.9081 (0.0248) 0.9252 (0.0123) 0.9274 (0.013) 0.9758 (0.0118) 0.8973 (0.0112)
RWSV-BL 0.9718 (0.0098) 0.9549 (0.0092) 0.9515 (0.0101) 0.972 (0.01) 0.9557 (0.0099) 0.96 (0.0097) 0.975 (0.0093) 0.9422 (0.0107)
ASV-HS 0.6221 (0.0435) 0.9106 (0.0265) 0.9094 (0.0263) 0.8678 (0.0468) 0.9194 (0.0456) 0.8721 (0.0582) 0.9275 (0.0295) 0.9569 (0.0168)
ASV-DHS 0.6426 (0.041) 0.9009 (0.0281) 0.8973 (0.0278) 0.8584 (0.045) 0.9186 (0.0387) 0.8765 (0.0519) 0.9264 (0.0296) 0.9583 (0.017)
ASV-HS-N 0.8701 (0.0577) 0.9388 (0.0206) 0.9371 (0.0205) 0.9014 (0.0335) 0.9317 (0.0374) 0.8996 (0.0448) 0.9498 (0.0245) 0.9614 (0.0156)
ASV-DHS-N 0.8753 (0.0539) 0.9355 (0.0217) 0.9329 (0.0219) 0.8988 (0.0314) 0.9338 (0.032) 0.9035 (0.0415) 0.9494 (0.0245) 0.963 (0.0155)
Table 2: Model performance under each data-generating process (DGP 1–8), evaluated using: MAE (mean absolute error), MCIW (mean 90% credible interval width), and EC (empirical 90% coverage). Standard deviations are shown in parentheses. For each DGP, the two models with the lowest average MAE and the four models with the narrowest average MCIW are highlighted in bold.
@@
Figure 3: Box plots of Mean Absolute Error (MAE) across 1,000 sample paths comparing Stochastic Volatility (SV), Markov-Switching Stochastic Volatility with 2 Regimes (MSSV2), Generalized AutoRegressive Conditional Heteroskedasticity (GARCH), Markov-Switching GARCH with 2 Regimes (MSGARCH2), Random Walk SV with Inverse Gamma Prior (RWSV), Random Walk SV with Bayesian LASSO (RWSV-BL), Adaptive SV with the horseshoe prior (ASV-HS), and ASV with dynamic horseshoe prior (ASV-DHS). Perfectly specified models, drawn in black, are SV and MSSV2 for DGP 1 and 2 and GARCH and MSGARCH2 for DGP 4 and 5. All models are misspecified for DGP 3 and 6.
As shown in Table 4.2 and Figure 3, the proposed ASV models consistently outperform or match the true models in DGPs with large volatility shifts, including DGPs 2, 3, 5, 6, and 8. In DGP 2, their MAE is only marginally higher than that of MSSV2, and in DGP 5, even slightly lower than that of MSGARCH2, both of which are correctly specified for those settings. Notable differences are found in DGPs 3, 6, and 8. In DGP 3, the average MAE of ASV models is 0.24 (sd = 0.07), compared to 0.46 (sd = 0.15) for the next best model, RWSV-BL, leading to 44% reduction. Similar differences are observed in DGPs 5 and 6. Such results underscore ASV’s ability in capturing abrupt structural changes.
@@
ASV models also perform reasonably well in scenarios where the underlying volatility is stationary or smoothly varying, as in DGPs 1, 4, and 7. SV and GARCH achieve the lowest average MAEs in DGPs 1 and 4 due to perfect specification; ASV models perform similarly to other misspecified models. In DGP 7, where volatility evolves smoothly without abrupt shifts, ASV outperforms all models including SV and RWSV. This suggests that although ASV is designed to capture abrupt structural changes, it still maintains competitive performance in settings that favor smooth or mean-reverting dynamics.
@@
Also, the simulation study highlights the limitations of HMS models, MSSV2 and MSGARCH2, which rely on a fixed number of latent regimes. These models perform poorly when the assumed regime structure does not align with the data. Notably, they perform worse than their single-regime counterparts SV and GARCH in scenarios with stationary volatility (DGPs 1 and 4). Their performance also deteriorates when the number discrete levels are more than 2 (DGPs 3, 6, and 8). Even in DGP 7, where volatility evolves smoothly without abrupt changes, the discrete regime-switching assumption introduces artificial segmentation, leading to subpar performance. In contrast, ASV models adaptively estimate volatility using global-local shrinkage, flexibly adjusting to a wide range of structural patterns without assuming a fixed number of regimes.
@@
The second and third row of Table 4.2 presents the ECs and the MCIWs of the 90% credible intervals of the eight Bayesian methods in the simulation study: SV, MSSV2, RWSV, RWSV-BL and 4 variants of ASV. ASV-HS and ASV-DHS achieve nominal 90% coverage in all DGPs except in DGPs 1 and 4. In DGP 1, they exhibit substantial undercoverage, with empirical coverage dropping to 63%, and in DGP 4, mild undercoverage at 86%. This is likely due to the models’ assumption of nonstationarity, which conflicts with the stationary assumption of the DGP. ASV-HS-N and ASV-DHS-N address the undercoverage problem by introducing a nugget effect, achieving correct nominal coverage across all 8 DGPs. In addition, their credible intervals are the narrowest across all eight DGPs aside from their non-nugget counter parts, indicating both precise and robust uncertainty quantification.
The key structural difference between ASV-DHS and ASV-HS is the presence of an autoregressive shrinkage parameter $\phi$ in ASV-DHS, which governs the temporal evolution of the local scale parameters. Despite this, they exhibit similar performance across DGPs 1 through 7. This is expected: in DGPs 1, 4, and 7, changes in volatility are gradual. In DGPs 2, 3, 5, and 6, abrupt changes occur but are highly persistent with the transition probability between different states of 2% for DGPs 2 and 5, and only 1% for DGPs 3 and 6, resulting in relatively few shifts. In DGP 8 features frequent and periodic volatility changes. In this setting, ASV-DHS consistently achieves smaller MAEs than ASV-HS. Although the average improvement is modest (0.021), it is consistent across 1,000 replicates, indicating that dynamic shrinkage offers better local adaptivity when volatility shifts are both abrupt and recurrent.

### v5

@@ 4.1 Set-up
The simulation study evaluates the performance of the proposed ASV model against several existing methods, including the SV, Markov-Switching SV with two regimes (MSSV2), RWSV, GARCH, and Markov-Switching GARCH with two regimes (MSGARCH2). A local-only shrinkage model, RWSV-BL, is also included for comparison. Four variants of ASV are considered: ASV with the horseshoe prior (ASV-HS), ASV with the dynamic horseshoe prior (ASV-DHS), and their respective versions with a nugget effect (ASV-HS-N and ASV-DHS-N). ASV-HS corresponds to the special case of Equation 4 with $a=b=1/2$ and $\phi=0$, while ASV-DHS generalizes this by allowing $\phi$ to be estimated from the data. All ASV models use $k=1$, applying shrinkage to the first-order differences of the latent log-variance process. This choice reflects the empirical observation that volatility in SV- and GARCH-type models typically exhibits clustering rather than gradual changes in slope.
Eight DGPs are considered: DGPs 1-3 are based on SV models with one to three regimes, while DGPs 4–6 are based on GARCH models with one to three regimes. To ensure that the simulated data reflect realistic volatility dynamics, the parameters for each DGP are calibrated using estimates obtained from empirical data. Specifically, DGPs 1 and 4 use parameters estimated from the S&P 500 data (Figure 1(a)), DGPs 2 and 5 from the electricity data (Figure 1(b)), and DGPs 3 and 6 from the bike rental data (Figure 1(c)). DGP 7 features sinusoidal log-variance with random coefficients, and DGP 8 features a piecewise constant log-variance with three to five random breakpoints. For each DGP, 1,000 sample paths are simulated. Detailed parameterization of the DGPs, including the transition probabilities, are provided in Appendix C.
@@
[…]
The MAE measures the accuracy of the volatility estimate by calculating the average absolute difference between the true volatility, $\sigma_{t}$, and the model’s estimate, $\hat{\sigma_{t}}$. For frequentist methods, such as GARCH and MSGARCH2, the point estimate corresponds to the maximum likelihood estimate, whereas for Bayesian methods, it corresponds to the posterior mean. The EC assesses the reliability of each model’s uncertainty quantification by calculating the proportion of true volatility values that fall within the model’s 90% credible interval. Finally, the MCIW evaluates the precision of this uncertainty quantification, with narrower credible intervals indicating higher precision provided that nominal coverage is achieved. Because GARCH and MSGARCH2 provide deterministic estimates of conditional volatility given the fitted parameters and data, MCIW and EC are not reported for these models.
All models are implemented in the R language (Rlang). Specifically, the SV model is implemented using the stochvol package (Stochvol); MSSV2, as specified in MSSVmine, and RWSV are implemented by the authors, as no readily available R packages exist. The GARCH(1,1) model is fitted using the fGarch package (fGarch), and MSGARCH2 is fitted using the MSGARCH package (MSGARCHP). To fit RWSV-BL, the dsp package by dsp and the genlasso package by genlasso are used. The dsp package is extended by the authors to implement ASV. For Bayesian models, 5,000 posterior samples are retained after 20,000 burn-in iterations.
@@ 4.2 Results
MAE SV 4.267e-03 (4.581e-04) 16.6293 (3.9431) 2.1691 (0.4344) 0.0245 (0.0055) 4.1234 (0.4709) 1.1816 (0.261) 1.1831 (0.9398) 2.0796 (0.1713)
MSSV2 4.302e-03 (4.781e-04) 6.7195 (1.9145) 2.0253 (0.3813) 0.0248 (0.0056) 3.0737 (0.7494) 1.0478 (0.2646) 1.0434 (0.8259) 1.4738 (0.2381)
GARCH 5.982e-03 (7.158e-04) 34.3626 (8.3057) 3.5958 (0.7481) 0.0047 (0.0026) 5.6249 (0.6527) 1.7023 (0.3995) 2.7572 (2.2338) 4.1509 (0.3525)
MSGARCH2 5.909e-03 (7.031e-04) 17.8737 (1.866) 2.8629 (0.5375) 0.0063 (0.0036) 3.8062 (0.6359) 1.2492 (0.3289) 2.1427 (1.6771) 2.7224 (0.7748)
RWSV 4.359e-03 (4.699e-04) 16.3886 (3.8574) 2.1386 (0.421) 0.0251 (0.0054) 4.0889 (0.4599) 1.1638 (0.254) 1.1668 (0.925) 2.0301 (0.1692)
RWSV-BL 5.223e-03 (5.582e-04) 19.4377 (4.6705) 2.5723 (0.5261) 0.0319 (0.0068) 5.742 (0.6862) 1.6914 (0.4265) 1.42 (1.0283) 1.9933 (0.1658)
[4pt/2pt] ASV-HS 4.562e-03 (4.918e-04) 13.0979 (2.836) 1.4824 (0.3262) 0.026 (0.0056) 3.4967 (0.5326) 0.9538 (0.23) 0.9348 (0.7365) 1.2409 (0.1392)
ASV-DHS 4.601e-03 (4.976e-04) 12.4654 (2.6887) 1.4365 (0.3183) 0.0262 (0.0056) 3.4395 (0.542) 0.9551 (0.2301) 0.9426 (0.7442) 1.2209 (0.1385)
ASV-HS-N 4.562e-03 (4.940e-04) 13.1351 (2.8391) 1.4866 (0.3266) 0.0262 (0.0056) 3.5331 (0.5424) 0.9603 (0.2311) 0.9358 (0.7373) 1.2449 (0.1386)
ASV-DHS-N 4.592e-03 (4.977e-04) 12.5345 (2.6921) 1.4475 (0.3189) 0.0263 (0.0056) 3.4831 (0.545) 0.9619 (0.2316) 0.9427 (0.7423) 1.2242 (0.1382)
MCIW SV 0.01787 (0.002083) 93.5561 (22.6665) 10.4085 (2.3493) 0.1055 (0.0207) 18.7136 (2.4446) 6.1095 (1.4139) 6.9224 (5.4048) 10.5937 (0.8248)
MSSV2 0.01745 (0.002244) 35.6871 (10.954) 9.3695 (2.015) 0.1003 (0.0202) 20.0305 (3.6871) 4.4687 (1.2165) 5.3407 (4.2339) 6.6075 (1.0614)
@@
RWSV 0.01527 (0.001879) 91.3216 (22.0039) 9.8339 (2.211) 0.0955 (0.0215) 16.5878 (2.0431) 5.8831 (1.3353) 6.801 (5.3228) 10.0901 (0.7914)
RWSV-BL 0.02885 (0.002798) 114.9675 (27.5597) 14.4477 (2.9798) 0.1808 (0.0331) 33.4065 (3.6979) 10.0222 (2.5165) 8.4838 (6.0041) 10.6267 (0.7472)
[4pt/2pt] ASV-HS 0.01415 (0.001847) 65.2994 (13.9958) 6.6685 (1.3651) 0.0903 (0.0217) 14.2277 (1.8101) 4.4709 (0.9961) 4.2954 (3.2816) 6.4435 (0.4201)
ASV-DHS 0.01396 (0.001849) 60.8098 (12.7929) 6.4654 (1.2943) 0.0892 (0.0216) 13.9008 (1.763) 4.4124 (0.9761) 4.3191 (3.3025) 6.3782 (0.4163)
ASV-HS-N 0.01736 (0.002667) 68.8245 (14.8542) 7.3104 (1.4842) 0.0948 (0.0221) 16.2665 (2.2526) 4.8062 (1.0979) 4.6034 (3.4542) 6.6306 (0.4301)
ASV-DHS-N 0.01742 (0.002744) 65.1302 (13.6983) 7.1792 (1.4312) 0.0943 (0.022) 16.0765 (2.1986) 4.7761 (1.0848) 4.6343 (3.4943) 6.5572 (0.4303)
EC SV 0.9037 (0.0276) 0.9584 (0.0125) 0.9134 (0.0162) 0.9323 (0.0136) 0.9056 (0.0251) 0.9322 (0.0152) 0.9764 (0.0114) 0.9091 (0.0107)
MSSV2 0.89 (0.04476) 0.9482 (0.0352) 0.9079 (0.0234) 0.9061 (0.036) 0.9618 (0.0522) 0.9059 (0.0457) 0.9543 (0.0251) 0.9539 (0.0237)
@@
RWSV 0.8326 (0.04226) 0.9555 (0.0132) 0.9002 (0.018) 0.8884 (0.021) 0.8665 (0.0314) 0.9261 (0.0168) 0.9758 (0.0118) 0.8973 (0.0112)
RWSV-BL 0.9637 (0.01033) 0.9696 (0.0094) 0.9593 (0.0108) 0.9675 (0.0101) 0.9685 (0.0096) 0.9686 (0.0097) 0.975 (0.0093) 0.9422 (0.0107)
[4pt/2pt] ASV-HS 0.7652 (0.04822) 0.9398 (0.0203) 0.9326 (0.0279) 0.8512 (0.0349) 0.9024 (0.0395) 0.9306 (0.0274) 0.9275 (0.0295) 0.9569 (0.0168)
ASV-DHS 0.7526 (0.04992) 0.9371 (0.0228) 0.9332 (0.0291) 0.8398 (0.0367) 0.9027 (0.0412) 0.9264 (0.029) 0.9264 (0.0296) 0.9583 (0.017)
ASV-HS-N 0.8587 (0.04874) 0.9489 (0.0178) 0.9502 (0.0231) 0.8711 (0.0291) 0.9269 (0.0315) 0.9452 (0.0229) 0.9498 (0.0245) 0.9614 (0.0156)
ASV-DHS-N 0.8567 (0.05088) 0.9493 (0.0193) 0.9531 (0.023) 0.8643 (0.0303) 0.9288 (0.032) 0.9431 (0.0238) 0.9494 (0.0245) 0.963 (0.0155)
Table 1: Model performance under each data-generating process (DGPs 1–8), evaluated using: MAE (mean absolute error), MCIW (mean 90% credible interval width), and EC (empirical 90% coverage). Standard deviations are shown in parentheses. For each DGP, the two models with the lowest average MAE, the model closest to the nominal 90% coverage, and the model with the narrowest average MCIW are highlighted in bold. MCIW and EC are not reported for GARCH and MSGARCH2, as these frequentist models yield deterministic estimates of conditional volatility given the fitted parameters and data.
@@
Figure 3: Box plots of mean absolute error (MAE) across 1,000 sample paths comparing Stochastic Volatility (SV), Markov-Switching Stochastic Volatility with 2 Regimes (MSSV2), Generalized Autoregressive Conditional Heteroskedasticity (GARCH), Markov-Switching GARCH with 2 Regimes (MSGARCH2), Random Walk SV with Inverse-Gamma Prior (RWSV), Random Walk SV with Bayesian LASSO (RWSV-BL), Adaptive Stochastic Volatility with the Horseshoe prior (ASV-HS), and Adaptive Stochastic Volatility with the Dynamic Horseshoe prior (ASV-DHS). Perfectly specified models, drawn in black, are SV and MSSV2 for DGPs 1 and 2, and GARCH and MSGARCH2 for DGPs 4 and 5. All models are misspecified for DGPs 3,6,7 and 8.
As shown in Table 4.2 and Figure 3, the proposed ASV models consistently outperform or match the true models in DGPs with large volatility shifts, including DGPs 2, 3, 5, 6, and 8. In DGPs 2 and 5, MSSV2 achieves the lowest MAE, as expected, since it assumes two distinct volatility states that match the structure of these DGPs. The proposed ASV models, despite not assuming a fixed two-state structure, consistently achieve the second-lowest MAE in both cases.
ASV models outperform existing alternatives particularly when large and abrupt changes in volatility are present. In DGP 3, the average MAE of the ASV models is 1.44 (sd = 0.32), compared with 2.03 (sd = 0.38) for the next best model, MSSV2, corresponding to a 29% reduction. Similarly, DGP 8 shows an 18% reduction in MAE, and DGP 6 shows a 9% reduction relative to the second-best model. Despite having three distinct states, existing models perform relatively well in DGP 6. This is primarily due to the high persistence parameter, $\beta$, characterizing DGP 6, which enables autoregressive models to track volatility changes more effectively. When persistence is lower, as in DGPs 3 and 8, the advantage of ASV becomes substantially more pronounced. These results highlight the strong ability of ASV to capture abrupt structural shifts in volatility dynamics.
@@
ASV models also perform reasonably well in scenarios where the underlying volatility is stationary or smoothly varying, as in DGPs 1 and 7. In DGP 1, the SV model achieves the lowest average MAEs, as it is the correctly specified model. The ASV models perform similarly to other misspecified SV type models, such as RWSV and MSSV2. In DGP 7, where volatility evolves smoothly without abrupt shifts, the ASV models outperform all alternatives, including SV and RWSV. This suggests that although ASV is designed to capture abrupt structural changes, it remains competitive in settings characterized by smooth or mean-reverting dynamics. ASV performs relatively poorly in DGP 4, showing MAE values comparable to other SV-type models. The reduced performance likely stems from model misspecification, as DGP 4 follows a GARCH(1,1) process while ASV assumes an SV formulation. GARCH volatility is a deterministic function of past data and therefore produces a much rougher stationary volatility path than the smoother dynamics in SV models. This mismatch makes accurate volatility estimation more difficult for ASV.
@@
The simulation study also highlights the limitations of regime-switching models such as MSSV2 and MSGARCH2, which rely on a fixed number of latent regimes. These models perform poorly when the assumed regime structure does not align with the true DGPs. This mismatch is evident in DGPs 3, 6, and 8, where the number of volatility levels exceeds the models’ assumed structure. In contrast, the ASV models estimate volatility adaptively through a global-local shrinkage prior, allowing flexible adjustment to a wide range of structural patterns without requiring a fixed regime specification.
@@
The second and third rows of Table 4.2 present the ECs and MCIWs of the 90% credible intervals for the eight Bayesian methods in the simulation study: SV, MSSV2, RWSV, RWSV-BL and four variants of ASV. They are not reported for GARCH and MSGARCH2, as these frequentist models yield deterministic estimates of conditional volatility given the fitted parameters and data. ASV-HS and ASV-DHS achieve near-nominal 90% coverage in all DGPs except DGPs 1 and 4. In DGP 1, they exhibit substantial undercoverage, with empirical coverage dropping to 76%, and in DGP 4, they show mild undercoverage at 84%. This is likely due to the models’ assumption of nonstationarity, which conflicts with the stationary structure of the DGPs 1 and 4. The nugget-augmented variants, ASV-HS-N and ASV-DHS-N, address this issue by introducing an additional noise term, achieving improved coverage of 86% in DGP 1 and 87% in DGP 4.
Across all DGPs, the ASV models produce narrow credible intervals, achieving the smallest MCIWs in seven of the eight settings. MSSV2 yields the narrowest intervals in DGP 2, where it is the correctly specified model. However, the undercoverage observed for ASV in DGP 1 suggests that its intervals are too narrow, leading to an underestimation of uncertainty. The ASV-N variants yield intervals that are only slightly wider than those of the base ASV models, yet remain generally narrower than those of other competing models. These results indicate that ASV models with a nugget term provide both precise and well-calibrated uncertainty quantification.
The main structural difference between the HS and DHS variants of ASV is the inclusion of the autoregressive shrinkage parameter, $\phi$, in ASV-DHS, which controls the temporal dependence of the local scale parameters. Compared with ASV-HS, ASV-DHS shows consistent but modest improvements in DGPs 2, 3, 5, and 8, where large and recurrent volatility shifts occur, with reductions in MAE of 5%, 3%, 2%, and 2%, respectively. These results suggest that dynamic shrinkage offers improved local adaptivity when volatility shifts are both abrupt and recurrent.